OpenStudy (anonymous):

1/2x +1/3 y = 3 Which of the following equations is equivalent to the given equation?

OpenStudy (anonymous):

which following

OpenStudy (anonymous):

3x + 2y = 3 2x + 3y = 18 3x + 2y = 18

OpenStudy (anonymous):

the last one is eqivalent